### v3

@@ Conclusion
We would like to emphasize that the human pose estimation tools are unfortunately not standard computer programs that can be easily deployed and used “as is”. Their installation may be more or less involved. Moreover, performance is affected by preprocessing steps and settings. For example, cropping images so that only the infant is in the image helps to improve performances. Images with multiple persons in the scene (e.g., caregiver or experimenter), especially in close proximity, are currently challenging, and body keypoints of adults and the infant can even get mixed up. Also, as the pose estimation networks were predominantly trained on people in upright positions, the orientation of the infant image importantly affects the performance—the camera should be positioned or every frame rotated afterwards so that the infant is positioned approximately with the head at the top. Some pose estimation networks may have additional settings (e.g., OpenPose and MediaPipe allow to set the maximum number of persons to detect in the image).
@@
Pose estimation methods can take either videos or individual images as input (or both). Although this may seem equivalent at first sight—a video is a sequence of consecutive images—, we realized that we sometimes obtained quite different results for these two input modes. The differences might partially come from how each method cuts the videos into frames, which can lead to slight offsets compared to the ground-truth sequences we used for evaluation. However, this alone cannot explain the larger differences in performance for OpenPose or the different in the number of redundant detections for Detectron 2. Some of the pose estimation methods have divergent processing paths for image and video processing. In particular, tracking, a step that exploits the temporal consistency of estimations over the sequence of detections, when implemented, is often enabled only for video input and not for images. With the exception of MediaPipe, for which we enabled it for both input modes in our ”supine” and MINI-RGBD datasets thanks to available parameters, none of the methods we tested had a clear tracking option in its parameters. Tracking can be currently achieved in postprocessing, as not all methods contain such a step, but in the future it is likely to find its way into the pose estimation methods themselves.
The top-down methods, which identify areas (bounding boxes) where persons are likely to be in the images first, and then look for keypoints on the body within these areas, require a (bounding-box) detector that provides the coordinates of this bounding box, in theory encompassing all the visible body parts of detected individuals. They are the main source of redundant and missed detections for these methods. Detectors are separate from the methods but often embedded within them, and some work might be needed to compare their performance in detecting infants. For example, we hypothesize that the poor performance of MediaPipe is influenced by its detector, which was made to detect (adult) faces instead of bodies and assumes that the body is directly below, as a compromise to increase processing speed. We can also observe that on our ”Lap” dataset, it is the bottom-up methods, HRNet BU and OpenPose that have the lowest Neck-MidHip errors, while all top-down methods show much higher errors. Another example is by looking at ViTPose’s reported results on the OCHuman dataset: a dataset comprised of multi-body scenes with adults. The AP for ViTPose-b is reported to be 87.3 when using the ground truth bounding boxes ([49] Table 11), while another work comparing the performance of several detectors on this dataset showed that the AP of ViTPose-b using the RTMDet detector drops to 45.0; and even the best detector only reached 48.3 AP on this dataset ([60], Table 1.
@@
Such results could be explained by the dataset containing infants only in supine position with a view from above, which, as long as the infants do not perform complex leg movements, is rather simple and could be considered similar enough to an adult standing in front of a camera: the situation that these methods encounter most commonly during their training. This is supported by looking at the average OKS values from individual videos (Supp. Materials ST. 6), where the performance of all methods decreases noticeably and is more variable for “hard” synthetic infants (IDs 10-12) compared to their performance on “easy” infants (IDs 1-4). Furthermore, using more complex datasets, such as SyRIP or our ”Lap” dataset shows large drops of performance at all levels. The ”Lap” condition, which would fall under a multi-body detection scenario, is challenging even when only adults are involved, partially due to the poor performance of detectors for top-down methods [60]. In our results, ViTPose-h performs 15 AP points lower in our ”Lap” dataset compared to ViTPose-b on the OCHuman dataset, which might highlight a bias toward detecting adults rather than infants when both are present in a scene together due to the training datasets only containing adults. It also scores about 40 points lower in AP on the SyRIP dataset compared to the common benchmark dataset for adults, COCO.
There are several directions for future work. First, although small infants (under seven months) differ significantly in their body proportions from adults and are thus strongly “out of distribution” data for the models in terms of shape estimation, pose estimation may be relatively easier because infants are in supine position all the time and their motor repertoire is constrained. Older infants and children in prone, standing, and other postures which are partially covered in the SyRIP dataset should be considered as a future challenge for pose estimation, and be included in the training of future methods considering the difficulty for current methods to handle such postures. The situation is similar for scenes with multiple people, like infants on a parent’s lap. Second, in addition to 2D pose estimation in the form of keypoints, there are methods that estimate 3D pose and shape (a complete body mesh) from an image or video (e.g., Smplify-x [58], with SMIL for an infant shape [61], or 4D Humans currently with adult shape only [62]). Third, while pose estimation is a key prerequisite for additional analyzes, the quality of motion extracted from the sequences of estimated keypoints should be studied in detail (see also [63]).
@@
Two of the datasets used (MINI-RGBD dataset [54] and SyRIP dataset [55]) are publicly available in their respective public repositories. Video recordings and images from our infant datasets (“supine” and “lap”) cannot be publicly shared due to data protection and sharing rules.
@@
In a best-case scenario (see Tab. S 4), pose estimation methods can reach high levels of performance on infants in supine position, with low variability, with ViTPose reaching an Average Precision up to 91.8, and average errors around 3.9$\pm$2.0of the Neck-MidHip segment (which corresponds roughly to the infant’s torso length).
With the detailed OKS values for each video (Tables ST. 5 and 6), we could get further insight with regards to the complexity estimation made by Hesse et al. on their MINI-RGBD dataset based on the actual performance of the methods’ estimations. Such a table can also help to identify for each specific method which kind of sequences they seem to struggle with to identify the possible causes of keypoint misplacement, so that more of such examples can be included in future methods’ training or fine-tuning. For example, MediaPipe did not manage to identify a single image among the sequence for synthetic infant 2, despite it being deemed ”easy”, possibly due to its unique background.
From Table 7 of the main manuscripts and Tables of Supplementary Materials ST 3 and ST 4, it seems that the better the estimations and the easier the videos, the lower the correlation between scores and OKS values. This could be explained by a lower range of high values from OKS, while the scores might be less reliable and might not reflect a similar reduction in their variability and range of values.

### v4

@@ Conclusion
We would like to emphasize that the human pose estimation tools are unfortunately not standard computer programs that can be easily deployed and used “as is”. Their installation may be more or less involved. Moreover, performance is affected by preprocessing steps and settings. For example, cropping images so that only the infant is in the image helps to improve performances. Images with multiple persons in the scene (e.g., caregiver or experimenter), especially in close proximity, are currently challenging, and body keypoints of adults and the infant can even get mixed up. Also, as the pose estimation networks were predominantly trained on people in upright positions, the orientation of the infant image importantly affects the performance—the camera should be positioned, or every frame rotated afterwards, so that the infant is positioned approximately with the head at the top. Some pose estimation networks may have additional settings (e.g., OpenPose and MediaPipe allow to set the maximum number of persons to detect in the image).
There are also differences between keypoint estimations within a single method. Some of these differences come from postural differences between infants and the datasets on which the methods were trained (featuring almost exclusively adults)—for example when legs are brought upwards, or are bent or crossed, as in Fig. 1, e). Other differences could reflect the accuracy of the human annotators (see the article and website for the COCO dataset [50] and how the $\sigma_{k}$ specific to each keypoint used for OKS were computed based on the variability between human annotators, as well as the differences we find between our two annotators). Errors can also occur naturally as an interaction between body posture and camera angle: in supine position and at younger ages, the head will have a tendency to be tilted in one direction rather than the other, which can lead to the occlusion of one ear that results either in a missing keypoint or in a less accurate estimation. Bottom-Up methods seem more prone to provide misestimations over the whole body when a subset of keypoints are suddenly misestimated or missing (e.g., the infant bringing its legs upwards and crossing the feet can lead to misestimation of the upper body). Our impression is that pose estimation methods seem to be able to estimate locations that are occluded but that could be guessed by a human annotator: for example, if the right arm is flexed and the right shoulder is hidden by the forearm or the hand, the shoulder keypoint will still be estimated with good enough accuracy.
We have observed issues when estimating images with higher luminance, on recordings where the infant was near windows during a sunny day. We have observed difficulties when the infant and the background are less contrasted, for example with white clothes on the infant with a white sheet under it, but we have also observed cases where the estimations looked accurate with infants having black clothes with a black sheet under (most likely because the cloth did not cover the lower legs and arms, and the infant’s skin color provided enough contrast). In other cases, we observed higher estimation errors when the camera angle was too low or too much on the side compared to a view from above. The distance between the camera and the infant and the zoom level also play a role [37], as well as the resolution (in particular if the infant has a pacifier).
@@
Pose estimation methods can take either videos or individual images as input (or both). Although this may seem equivalent at first sight—a video is a sequence of consecutive images—, we realized that we sometimes obtained quite different results for these two input modes. The differences might partially come from how each method cuts the videos into frames, which can lead to slight offsets compared to the ground-truth sequences we used for evaluation. However, this alone cannot explain the larger differences in performance for OpenPose or the difference in the number of redundant detections for Detectron 2. Some of the pose estimation methods have divergent processing paths for image and video processing. In particular, tracking, a step that exploits the temporal consistency of estimate over the sequence of detections, when implemented, is often enabled only for video input and not for images. With the exception of MediaPipe, for which we enabled it for both input modes in the “Supine” and MINI-RGBD datasets, none of the methods we tested had a clear tracking option in its parameters. Tracking can be currently achieved in postprocessing, as not all methods contain such a step, but in the future it is likely to find its way into the pose estimation methods themselves.
The top-down methods, which identify areas (bounding boxes) where persons are likely to be in the images first, and then look for keypoints on the body within these areas, require a (bounding-box) detector that provides the coordinates of this bounding box, in theory encompassing all the visible body parts of detected individuals. They are the main source of redundant and missing detections for these methods. Detectors are separate from the methods but often embedded within them, and some work might be needed to compare their performance in detecting infants. For example, we hypothesize that the poor performance of MediaPipe is influenced by its detector, which was made to detect (adult) faces instead of bodies and assumes that the body is directly below, as a compromise to increase processing speed. We can also observe that on our “Lap” dataset, it is the bottom-up methods, HRNet BU and OpenPose that have the lowest Neck-MidHip errors, while all top-down methods show much higher errors. Another example is by looking at ViTPose’s reported results on the OCHuman dataset: a dataset comprised of multi-body scenes with adults. The AP for ViTPose-b is reported to be 87.3 when using the ground truth bounding boxes ([49] Table 11), while another work comparing the performance of several detectors on this dataset showed that the AP of ViTPose-b using the RTMDet detector drops to 45.0; and even the best detector only reached 48.3 AP on this dataset ([61], Table 1).
@@
Such results could be explained by the dataset containing infants only in supine position with a view from above, which, as long as the infants do not perform complex leg movements, is rather simple and could be considered similar enough to an adult standing in front of a camera: the situation that these methods encounter most commonly during their training. This is supported by looking at the average OKS values from individual videos (Supp. Materials ST. 6), where the performance of all methods decreases noticeably and is more variable for “hard” synthetic infants (IDs 10-12) compared to their performance on “easy” infants (IDs 1-4). Furthermore, using more complex datasets, such as SyRIP or our “Lap” dataset, shows large drops of performance at all levels. The “Lap” condition, which would fall under a multi-body detection scenario, is challenging even when only adults are involved, partially due to the poor performance of detectors for top-down methods [61]. In our results, ViTPose-h performs 15 AP points lower in our “Lap” dataset compared to ViTPose-b on the OCHuman dataset, which might highlight a bias toward detecting adults rather than infants when both are present in a scene together due to the training datasets only containing adults. It also scores about 40 points lower in AP on the SyRIP dataset compared to the common benchmark dataset for adults, COCO.
There are several directions for future work. First, although small infants (under seven months) differ significantly in their body proportions from adults and are thus strongly “out of distribution” data for the models in terms of shape estimation, pose estimation may be relatively easier because infants are in supine position all the time and their motor repertoire is constrained. Older infants and children in prone, sitting, standing, and other postures which are partially covered in the SyRIP and our “Lap” datasets should be considered as a future challenge for pose estimation, and be included in the training of future methods considering the difficulty for current methods to handle such postures. The situation is similar for scenes with multiple people, like infants on a parent’s lap. Second, in addition to 2D pose estimation in the form of keypoints, there are methods that estimate 3D pose and shape (a complete body mesh) from an image or video (e.g., Smplify-x [58], with SMIL for an infant shape [62], or 4D Humans currently with adult shape only [63]). Third, while pose estimation is a key prerequisite for additional analyzes, the quality of motion extracted from the sequences of estimated keypoints should be studied in detail (see also [64]).
Pose estimation methods have been used in clinical movement assessment of infants (OpenPose outputs used for automated GMA scoring in [30] and for ASD prediction in [33]; AlphaPose used in [20] as input for subsequent classification using HINE). Here, we show that with ViTPose, the errors are considerably smaller (about 25% reduced magnitude of Neck-MidHip errors compared to OpenPose and AlphaPose for infants in supine position, and little to no missing data). Therefore, using state-of-the-art methods, together with the code and additional recommendations on the processing we provide here, holds promise for improving automated infant clinical assessment in the very near future.
@@
Two of the datasets used (MINI-RGBD [54] and SyRIP [55]) are publicly available in their respective public repositories. Video recordings and images from our infant datasets (“Supine” and “Lap”) cannot be publicly shared due to data protection and sharing rules.
@@
[…]
With the detailed OKS values for each video (Tables ST. 6 and 7), we could get further insight with regards to the complexity estimation made by Hesse et al. on their MINI-RGBD dataset based on the actual performance of the methods’ estimates. Such a table can also help to identify for each specific method which kind of sequences they seem to struggle with to identify the possible causes of keypoint misplacement, so that more of such examples can be included in future methods’ training or fine-tuning. For example, MediaPipe did not manage to identify a single image among the sequence for synthetic infant 2, despite it being deemed ”easy”, possibly due to its unique background.
From Table 7 of the main manuscripts and Tables of Supplementary Materials ST 4 and ST 5, it seems that the better the estimates and the easier the videos, the lower the correlation between scores and OKS values. This could be explained by a lower range of high values from OKS, while the scores might be less reliable and might not reflect a similar reduction in their variability and range of values.